If paving bricks are $45.50 per square metre, how much would it cost to pave a patio of 9.6m in length and 6.25m in width?
1
Expert's answer
2011-03-03T08:29:19-0500
Let's find the area of the patio:
S = 9.6*6.25 = 60 m2
Thus it would cost
60[m2] * 45.5[1/m2] = $2730
to pave it.